Find the Taylor series solution of following

(X2+1)y''+4xy'+(x-5)y=0 y(2)=-1

y'(2)=3


1
Expert's answer
2021-08-12T04:52:03-0400

Solution;

By definition, Taylor's series expansion of f(x) at x=a is given by;

"f(x)=f(a)+f'(a)(x-a)+\frac{f"(a)}{2!}(x-a)^2+\\frac{f'''(a)}{3!}(x-a)^3+..."

We are given X=2,hence;

"f(x)=f(2)+f'(2)(x-2)+\frac{f''(2)}{2!}(x-2)^2+\frac{f'''(2)}{3!}(x-2)^3+..."

Since we have;

y(2)=-1

y'(2)=3

Also;

(x2+1)y''+4xy'+(x-5)y=0

By substitution;

(22+1)y''+4(2)(3)+(2-5)(-1)=0

5y''=-27

y''(2)="\frac{-27}{5}"

Now ,find the triple derivative at x=2;

(x2+1)y''+4xy'+(x-5)y=0

Distribute;

"x^2\frac{d^2y}{dx}+\frac{d^2y}{dx^2}+4x\frac{dy}{dx}+xy-5y=0"

Differentiate;

"[x^2\frac{d^3y}{dx^3}+\frac{d^2y}{dx^2}2x]+\frac{d^3y}{dx^3}+[4x\frac{d^2y}{dx^2}+4\frac{dy}{dx}]+[x\frac{dy}{dx}+y]-0=0"

Rewrite as;

"[x^2y'''+2xy'']+y'''+[4xy''+4y']+[xy'+y]=0"

Substitute all known values;

"4y'''-\frac{27}{5}(4)+y'''-\frac{27}{5}(8)+4(3)+2(3)-1=0"

Simplify;

"5y'''-\frac{324}{5}+17=0"

Equate to obtain;

y'''(2)="\frac{239}{25}"

Hence;

"f(x)=-1+3(x-2)-\frac{27}{10}(x-2)^2+\frac{239}{150}(x-2)^3+..."

Distribute;"f(x)=-1+3(x-2)-\frac{27}{10}(x^2-4x+4)+\frac{239}{150}(x^3-6x^2+12x-8)+..."

Simplify;

"f(x)=-\frac{2291}{75}+\frac{823}{25}x-\frac{613}{50}x^2+\frac{239}{150}x^3+..."